### v1

# A Modular Safety Filter for Safety-Certified Cyber-Physical Systems

Abstract
Nowadays, many control systems are networked and embed communication and computation capabilities. Such control architectures are prone to cyber attacks on the cyberinfrastructure. Consequently, there is an impellent need to develop solutions to preserve the plant’s safety against potential attacks. To ensure safety, this paper introduces a modular safety filter approach that is effective for a variety of cyber-attack types. This solution can be implemented in combination with existing control and detection algorithms, effectively separating safety from performance. The safety filter does not require information on the reliability of the received command or the feature of the used anomaly detector. It can be implemented in conjunction with high-performance, resilient controllers, to achieve both high performance during normal operation and safety during an attack. As an illustrative example, we have shown the effectiveness of the proposed design considering a multi-agent formation task involving 20 mobile robots. The simulation results testify that the safety filter operates effectively during false data injection and intelligent attacks.
I INTRODUCTION
Cyber-Physical Systems (CPS) are networked control systems with a tight integration with computation and communication capabilities[1]. Merging cyber technologies with physical systems greatly boosts operational efficiencies. However, it also introduces vulnerabilities that undermine the reliability of essential infrastructure as the communication lines present opportunities for hackers to manipulate data lines and initiate cyber attacks. A variety of solutions have been proposed to prevent, detect, and mitigate cyber-attacks using control theoretical tools[2]. Most of the solutions consider systems without constraints and/or the mitigation strategies rely on the use of an anomaly detector to ensure that the system does not enter unsafe configurations.
Recently, there has been an increasing trend in constrained CPS to address safety concerns explicitly. This involves formally defining safe zones as constraints within state and input spaces. In this paper, a Modular Safety Filter (MSF), inspired by safety-certified learning-based controllers [3]-[4], is proposed to satisfy safety constraints in the presence of cyber-attacks on the actuator and sensor signals. Due to its modularity, this method can be used as a standalone technology alongside other resilient controllers and anomaly detectors. Performance and safety criteria are separated in our architecture, simplifying the design process. Since MSF makes no assumptions about the attacked signal, such as limited bandwidth and small bounds, or the attacker’s computational power, it can be used in a wide range of situations.
Prior work largely focuses on linear systems. In [5] and [6], (set-theoretic) MPCs are employed for linear systems under False Data Injection (FDI) attacks, guaranteeing the stability and constraint satisfaction of the system. In [7], a distributed MPC and attack detection framework is proposed for constrained linear multi-agent systems under adversarial attacks. In [8], a method for tracking problems that require reachable sets is proposed for constrained linear systems under arbitrary attacks on both the actuation and measurement lines. A data-driven approach for LTI systems is proposed in [9] and [10], introducing a safety verification plus emergency control module assuming only noise-polluted input-state trajectories are available. A semi-definite approach, assuming bounded additive attacker’s signals, is proposed in [11] to design a safety-preserving filter for deterministic LTI systems under FDI attacks. In [12], a set-theoretic receding horizon control has been proposed to address FDI and denial of service attacks for LTI systems. In [13] and [14], reachability analysis is used and investigated to design safety-preserving platforms for LTI systems. A solution for nonlinear systems is proposed in [15] and [16] which provides safety based on an invariant set of SOS-based Lyapunov functions, resulting in conservative ellipsoidal safe sets.
The majority of proposed methods are developed resorting to reachable set arguments. This reliance often leads to a focus on LTI systems, since they become hard-to-compute or expensive for high-order and nonlinear systems. There is an emerging trend to separate the control performance from the safety. However, defining performance and safety through a single objective may degrade performance and result in conservative solutions where the safety objective is not directly linked to the control objective. To achieve superior performance, one may aim to optimize a non-continuous cost function via a deep policy while ensuring safety constraints in the presence of an attack, a task not feasible with current solutions. The proposed architecture aims to address these gaps.
Figure 1: Proposed safety-certified architecture for cyber-physical systems.
The remainder of the paper is organized as follows. Preliminary material, adversarial capabilities, and the problem statement are described in section II. Section III includes the proposed MSF. A multi-agent setup consisting of 20 mobile robots is described in section IV. Simulation results obtained considering different FDI attacks are presented in section V. Lastly, conclusions and limitations of the present work are discussed in section VI.
II Problem Statement and Preliminary Material
This section provides an overview of the CPS problem considered in this paper and the proposed MSF architecture, depicted in Fig. 1. It also details the formulation of the physical system and the types of cyber-attacks used for simulation examples.
II-A Physical System
Consider a class of dynamical systems that can be described by a set of non-linear equations as follows,
$\dot{x}(t)=f(x(t),u(t)),$ (1)
where $t\in\mathbb{R}$ is the continuous time, $x=x(t)\in\mathbb{R}^{n}$ is the state vector, and $u=u(t)\in\mathbb{R}^{m}$ is the control input. For a unique solution, $f(.)$ is assumed to be continuously differentiable, and $u$ is a piecewise-continuous signal.
Definition 1 (Admissible Set).
An admissible set for the dynamical system (1) is defined as a subset of the state- and input-space as follows:
$\begin{split}u\in\mathcal{U}\subseteq\mathbb{R}^{m},x\in\mathcal{X}\subseteq\mathbb{R}^{n}.\end{split}$ (2)
Definition 2 (Safe Set).
A safe set, $\mathcal{S}\subseteq\mathcal{X}$, is a set of initial conditions for which there exists an input-state trajectory subjected to the dynamical system (1) that belongs to the admissible set (2) for an infinite time.
$\mathcal{S}=\{x(t_{0})\in\mathcal{X}\mid\exists u(t)\in\mathcal{U},x(t)\in\mathcal{X},\forall t>t_{0}\}.$ (3)
II-B Adversarial Capabilities
Let us assume that $x(t)$ is sent from the local network to the cyber layer via an unsecured network. The cyber layer computes the control action, $u_{c}=u_{c}(t)$, and transmits it back to the local network. These signals are susceptible to cyber-attacks, denoted by $x_{a}=x_{a}(t)$ and $u_{a}=u_{a}(t)$, respectively. Note that this scenario is similar to the scenarios considered in [9] and [11]. Without losing generality, a cyber attack can be described using the following unknown function:
$(u_{a},x_{a})=h(u_{c},x).$ (4)
Function $h(\cdot,\cdot)$ can be defined to represent different types of attack and adversarial capabilities [^1] The attacker may use previous inputs and states. Due to the simplicity of the notation, it is omitted here.. While the proposed modular safety filter does not rely on assumptions on the attack, i.e., the function $h(\cdot,\cdot)$ is unknown to the proposed safety filter, we will consider three situations in the illustrative simulation example in Section V: attack-free, FDI, and intelligent attacks.
II-B1 Attack-free
$h(\cdot,\cdot)$ is defined as the identity function in the absence of the attack as follows:
$u_{a}=u_{c},\ x_{a}=x.$ (5)
II-B2 FDI Attack
The attacker has the capability of adding arbitrary perturbations $\delta_{x}$ and $\delta_{u}$ to the control input and state measurement vectors.
$u_{a}=u_{c}+\delta_{u},\ x_{a}=x+\delta_{x}.$ (6)
II-B3 Intelligent Attack
The attacker is assumed to know the system dynamics and safe set, $\mathcal{S}$, as well as disclosure and disruptive resources on $x(t)$ and $u_{c}(t)$. Furthermore, the attacker has sufficient computational power to compute $h(\cdot,\cdot)$ resorting to any desired optimal policy. This includes an intelligent undetectable attack strategy that cannot be identified by an anomaly detector in the cyber layer. In general, a generic attack action can be modeled as the following optimization problem.
$\max\quad J(x,u_{c})$ (7)
[…]
where $J(\cdot,\cdot)$ is the attacker’s objective and the attacker’s constraints represents conditions to ensure the attack is undetectable by an anomaly detector. Such conditions depend on the used anomaly detector and are left unspecified for the sake of generality.
II-C Central Controller
The central control policy may be designed for regulation, tracking, or any other task.
$u_{c}=g(r,x_{a}),$ (8)
where $r$ represents the reference trajectory. It is assumed that the controller is safety-certified in the absence of FDIs. The function $g(\cdot,\cdot)$ is unknown to the proposed safety filter [^2] The central controller may use future reference and past input-state data. Due to the simplicity of the notation, it is omitted here..
III Modular Safety Filter
We propose an architecture for which a modular safety filter can be implemented as an independent add-on module by filtering the control signal. Upon receiving $x(t)$ and $u_{a}(t)$, the safety filter provides the nearest safe input, $u_{s}(t)$, to the command signal $u_{a}(t)$, while respecting the system constraints (2) at all times. The safety filter can be described by the following optimization problem,
$u_{s}$ $:=\operatorname*{arg\,min}_{u}\|u_{a}-u\|^{2}_{2}$ (9)
s.t. safety $\quad\text{\emph{safety constraints}},\quad\forall t>0$
where ||.$||.||_{2}$ is the 2-norm of a vector. The safety constraints represents satisfying condition (2). To solve this problem, inspired by [17], we propose a predictive-based safety filter to handle safety constraints at all times including when potentially unsafe inputs are presented. The predictive safety filter approximates the problem (9) by searching for a safe backup trajectory toward a final safe set with finite-time prediction. Consider a discrete version of (1) as follows,
$x(k+1)=f_{d}(x(k),u_{s}(k)),$ (10)
where $u_{s}(k)=u_{0}^{k}$ is the first element of the solution of the optimization (III). The discrete-time index is defined as $k=\frac{t}{T_{s}}$ with $T_{s}$ representing the sampling time. The predictive safety filter is outlined below,
$\min_{u_{n\in\mathcal{N}}^{k}}\|u_{a}(k)-u_{0}^{k}\|^{2}_{2}$ (11a)
[…]
$x_{n+1}^{k}=f_{d}(x_{n}^{k},u_{n}^{k}),$ (11c)
$(x_{n}^{k},u_{n}^{k})\in(\mathcal{X},\mathcal{U}),$ (11d)
$x_{N}^{k}\in\mathcal{S}_{f},$ (11e)
$x_{0}^{k}=x(k),$ (11f)
where (11a) yields the nearest safe action to $u_{a}(k)$, (11c) is the prediction model, (11d) is the admissible set, (11e) is the final safe set, (11f) is the initial condition at discrete time $k$, $u_{n}^{k}$ is the $n^{th}$ element of prediction at discrete-time index $k$, and $N$ is the prediction horizon. Note that if $u_{a}$ is safe, the safety filter does not alter the input. The solution to this optimization problem yields an input-state backup trajectory. The safe set $\mathcal{S}_{f}$ is defined to guarantee the recursive feasibility of the filter similar to the terminal condition in MPC [18]. A visual illustration of the admissible set, safe set, final safe set, and the backup trajectory at time $k$ is shown in Fig. 2. The safety filter algorithm is summarized in the Algorithm 1.
Assumption 1.
(Initial Feasibility) The optimization problem (III) is feasible at $k=0$.
Assumption 2.
(Final Safe Set, $\mathcal{S}_{f}$) Let $\mathcal{S}_{f}\subseteq\mathcal{X}$, and $\mathcal{K}_{f}$ be a local policy that renders system (10) invariant on $\mathcal{S}_{f}$.
Note that the safe set, $\mathcal{S}$, is implicitly considered via the MSF optimization problem (III). The size of this set depends on the prediction horizon and the size of the final safe set, $\mathcal{S}_{f}$. A short prediction horizon and a small final safe set may lead to an overly conservative solution. Generally, there is no need for a long prediction horizon to achieve a non-conservative solution, in contrast to a solution that aims to provide a robust performance controller as well as safety.
Lemma 1 (Proof of Safety).
Let Assumptions 1-2 hold. Then, the system (10) is safe in the sense of Definition 2.
Proof.
The proof follows from Theorem 4.6 in [17]. ∎
Note that safety remains ensured if the (non-convex) optimization is infeasible for some $k>0$, as a safe backup trajectory is available [17].
Figure 2: At the time $k$, an unsafe control input, $u_{a}(k)$, is received by the safety filter. Since applying this input may result in an unsafe trajectory (I) in the next steps, MSF will find a backup trajectory (III) towards the final safe set, $\mathcal{S}_{f}$, by applying the safe input, $u_{s}(k)$. Applying this safe input results in the modified trajectory (II).
Algorithm 1 Predictive Safety Filter for CPS
1:Initialize $\mathcal{S}_{f}$, $\mathcal{X}$, $\mathcal{U}$, $N$, $x(0)$, $k=0$.
2:while true do
3:     Solve problem (III) for $u_{a}(k)$.
4:     Apply $u_{s}(k)$ to system (1).
5:     Measure system’s states, $x(k+1)$, send it to the cyber layer and update the initial condition, $x_{0}^{k+1}$.
6:     $k\xrightarrow{}k+1$
7:end while
Remark: In this paper, we assumed that the underlying problem is deterministic, nominal, and has zero transmission delay. For other settings when a simplified model, probabilistic model, or additive disturbance is present, an adaptation of the algorithm is required which may introduce conservatism, see [3] and[4] for a recent overview of safety filter technology.
IV Centralized Predictive Safety Filter for Multi-Agent Mobile Robot System
To evaluate the efficiency of the proposed method on a high-order nonlinear system, we adopt the simulation framework employed in [19], which considered 20 linear mobile robots, extended with nonlinear dynamics as described in [20]. For simplicity, we assume that all mobile robots have the same dynamics and parameters. The kinematics model for the $i^{th}\in\mathcal{I}$ robot is described as follows:
$\dot{x_{i}}=v_{i}\cos\theta_{i},\quad\dot{y_{i}}=v_{i}\sin\theta_{i},\quad\dot{\theta_{i}}=\omega_{i},$ (12)
where $x_{i}$ and $y_{i}$ represents the position vector $p_{i}=[x_{i},y_{i}]^{\top}$, $\theta_{i}$ is the heading, $v_{i}$ and $\omega_{i}$ are the control inputs $u_{i}=[v_{i},\omega_{i}]^{\top}$, and =$\mathcal{I}=\{1,2,...,20\}$ is an index set indicating each agent. This non-holonomic model makes the control problem more challenging than a linear double integrator.
Figure 3: Schematic of mobile robots: linear and angular velocities $(v,\omega)$, and Cartesian coordinates $(x,y)$. The pre-defined safety constraints for the multi-agent system are the distance between two arbitrary robots $d_{i,j}$ and the distance between an arbitrary robot and a wall $d_{w_{i,j}}$.
We define the safety constraints through the admissible set (Definition 1), as the minimum distance between two arbitrary agents ($d_{[i,j]}\geq\delta_{a}$) and the minimum distance between one arbitrary agent and walls ($d_{w_{[i,j]}}\geq\delta_{w}$), described by (IV). A graphical visualization of these constraints for two agents as well as the inertial frame is depicted in Fig. 3. Note that constraint set (IV) is user-defined and only specifies the admissible set. The safe set is also dependent on the velocities, and actuator limitations which are implicitly considered by the MSF.
$d_{[i,j]}:=\{||p_{i}-p_{j}||_{2}:\forall i,j\in\mathcal{I},i\neq j\},\$ (13a)
$d_{w_{[i,j]}}:=\{||p_{i}-w_{j}||_{2}:\forall i\in\mathcal{I},\forall j\in\mathcal{J}\},$ (13b)
where, $w_{j}$ is the position of the $j^{th}\in\mathcal{J}:=\{1,2,3,4\}$ wall, $d_{[i,j]}$ represents the distance between $i^{th}$ and $j^{th}$ agents, $d_{w_{[i,j]}}$ represents the distance between $i^{th}$ agent and $j^{th}$ wall.
Note that the admissible set is explicitly defined via (IV), and the safe set is defined implicitly via the optimization problem (III), however, the final safe set still must be explicitly defined. For a multi-agent mobile system, we define the final safe set as a rest point for each robot when the velocity is equal to zero and the distance between two arbitrary robots is greater than a positive threshold (15d-15h). To avoid the collision, we define constraints over the backup trajectories for each agent as a minimum distance between the backup trajectories over the prediction horizon. The safety filter must be able to find safe backup trajectories that do not collide and have zero velocity at the end of the prediction horizon. We emphasize that this design is not case-dependent for mobile robots, it can be applied to similar scenarios like a group of aerial robots or any multi-agent system that has an equilibrium point. A circular reference trajectory with constant radius, $r_{0}$, and constant angular velocity, $\omega_{0}$, is defined as the formation task for the multi-agent mobile robot system as follows:
$\begin{array}[]{l}x_{i}^{d}=r_{0}\sin(w_{0}t+\frac{2\pi}{|\mathcal{I}|}(i-1)),\\ y_{i}^{d}=r_{0}\cos(w_{0}t+\frac{2\pi}{|\mathcal{I}|}(i-1)),\\ \end{array}$ (14)
where $|\mathcal{I}|$ is the cardinality of $\mathcal{I}$. The centralized safety filter for the multi-agent mobile robot system is defined as follows:
$U_{s}=\operatorname*{arg\,min}_{U_{n\in\mathcal{N}}^{k}}||U_{a}-U_{0}^{k}||^{2}_{2}$ (15a)
[…]
$X_{n+1}^{k}=f_{d}(X_{n}^{k},U_{n}^{k}),$ (15c)
$U_{n}^{k}\in\mathcal{U},$ (15d)
$X_{0}^{k}=X(k),$ (15e)
${d_{n}^{k}}_{[i,j]}\leq\delta_{a},\forall i\in\mathcal{I},\quad\forall j\in\mathcal{J},\quad\forall n\in\mathcal{N},$ (15f)
${{d_{w}}_{n}^{k}}_{[i,j]}\leq\delta_{w},\forall i\in\mathcal{I},\quad\forall j\in\mathcal{J},\quad\forall n\in\mathcal{N},$ (15g)
${d_{N}^{k}}_{[i,j]}\leq\delta_{a},\quad{{d_{w}}_{N}^{k}}_{[i,j]}\leq\delta_{w},\quad{v}_{N}^{k}=0,$ (15h)
where ${X}$ and ${U}$ are the stacked states and control inputs for all agents defined as $X=[x_{1},x_{2},...,x_{20}]^{\top}$ and $U=[u_{1},u_{2},...,u_{20}]^{\top}$, respectively. ${d_{n}^{k}}_{[i,j]}$ and ${{d_{w}}_{n}^{k}}_{[i,j]}$ are the distance similar to (IV) at time $k$ and $n^{th}$ prediction element. Equations (15f-15h) are defined for the safe backup trajectory to avoid collisions.
V Numerical Results
Two attack scenarios are implemented, following the attack definition in section II-B and safety filter setup in section IV. Note that the settings for the safety filter are identical for both scenarios as the safety filter does not depend on the type of attack. A description of the simulation setup and safety filter can be found in Table I. This simulation takes $20$ seconds, and attacks are applied in $t\in[5,10]$. We also employed an MPC controller for the tracking problem with a predictive horizon equal to 100, representing function $g(\cdot,\cdot)$ in section II-C, respectively. Note that the prediction horizon for the safety filter is $N=3$.
TABLE I: Simulation Parameters
Centralized Model-Predictive Safety Filter
Parameter Value Parameter Value
$N$ 3 $T_{s}^{\text{filter}}$ $0.02$ [sec]
$\delta_{a}$ $0.2$ [m] $\delta_{w}$ $0.2$ [m]
Multi-Agent Mobile Robot System
Parameter Value Parameter Value
$w_{0}$ $0.1\,[\frac{\text{rad}}{\text{s}}]$ $|\mathcal{I}|$ 20
[…]
$v_{\text{min}}$ $-2\,[\frac{m}{s}]$ $v_{\text{max}}$ $+2\,[\frac{m}{s}]$
$\omega_{{\text{min}}}$ $-2\,[\frac{\text{rad}}{\text{s}}]$ $\omega_{{\text{max}}}$ $+2\,[\frac{\text{rad}}{\text{s}}]$
[…]
V-A Intelligent Attack
Let the attacker read and manipulate the control and sensor measurement signals to perform an undetectable covert attack [21], for $k\in[500,1000]$. The attack vector on the measurement signals is described as follows:
$x_{a}(k+1)=f_{d}(x_{a}(k),u_{c}(k)),$ (16)
where $x_{a}$ at $k=500$, is equal to the system’s state. The evolution of equation (16) provides a state trajectory that an anomaly detector expects to see in the cyber layer, based on the control signal provided by the central controller, $u_{c}$. The attack vector, $u_{a}$, is computed via an optimization problem whose objective is to cause a collision at the origin. The results of simulation using the CasADi toolbox [22] for three snapshots at $t=0.1[s]$, $t=6.5[s]$, $t=20[s]$ representing before, during, and after the attack period are shown in Fig.(4-6). The regular system, shown on the left, has no safety shield while the safety-certified system, shown on the right, uses the filter defined via Algorithm (1). A video of this simulation is presented here [^3] Intelligent Attack: https://www.youtube.com/watch?v=cprja-LznkI.
As depicted in Fig. 4, the safety filter has no impact on the formation task before the attack, and agents converge to the circular trajectory (14), shown by the green circle, from their initial conditions. Fig. 5 illustrates the impact of the safety filter during the attack, which maintains the system’s safety and prevents collisions. Finally, Fig. 6 demonstrates how the system can recover itself once the attack is finished.
Figure 4: The multi-agent mobile robot system assigned to a formation task: following a circular trajectory at $t=0.1[s]$. (Before the intelligent attack)
Figure 5: The multi-agent mobile robot system assigned to a formation task: following a circular trajectory at $t=6.5[s]$. (During the intelligent attack)
Figure 6: The multi-agent mobile robot system assigned to a formation task: following a circular trajectory at $t=20[s]$. (After the intelligent attack)
Figure 7: The effect of the proposed modular safety filter on the first agent’s input. (Intelligent attack)
Fig. 7 displays the safety-certified and attack inputs for the first agent. For the range $t\in[0,5.49]\cup[10,20]$, the first agent remains safe since there is no modification to the input. It is important to note that for $t\in[5,5.49]$, demonstrated by the magenta area, attack input is applied, however, there is no modification as the corresponding agent remains safe. For $t\in[5.49,10]$, as depicted by the yellow area, there is a significant correction by the filter to prevent unsafe situations. During $t\in[10,20]$, after the attack period, the system successfully recovers to its normal condition, and the safety filter has zero impact.
V-B False Data Injection Attack
For all agents, we consider an FDI attack where $(v,\omega)=(v_{max},0)$. The objective of this attack is to steer all the agents outside of the admissible set, $\mathcal{X}$, as shown by the black square. The results are shown in Fig. 8 at $t=5.49[s]$. The safety-certified system prevents the agents from leaving the admissible set by forcing them to stop before reaching the boundaries. Since the results of non-attacked periods are similar to those of intelligent attacks, they are not illustrated in this paper, please visit the Link [^4] FDI Attack: https://www.youtube.com/watch?v=kBO05D3sZiE.
Figure 8: The multi-agent mobile robot system assigned to a formation task: following a circular trajectory at $t=5.49[s]$. (During the FDI attack)
VI Conclusions and Future Works
This paper presents a modular safety filter for Cyber-Physical Systems designed to ensure safety at all times, including in the presence of attacks. The safety filter ensures safety regardless of whether the received control command is safe or compromised by an attack, making it effective across various attacks without any assumptions on the attack model. Demonstrating the separation of safety and performance criteria, the proposed solution allows for safety during attacks alongside any high-performance controller. This highlights the versatility of safety filters in CPS applications, especially given that constrained controllers like MPC cannot optimize all types of cost functions.
The proposed safety filter is inspired by predictive safety filters developed for learning control. This paper illustrates the effectiveness of a modular approach to the safety of CPS that can handle nonlinear and high-order systems. Depending on the characteristics of the system, alternative safety filter solutions proposed for learning control can likely be used with minor adjustments to CPS. These methods include Control Barrier Functions (CBFs) and Hamilton-Jacobi (HJ) analysis where their extensions can account for uncertain, time-delay, and stochastic settings [23, 24, 25]. Each method comes with its advantages and disadvantages, but it is worth noting that the calculation of safe sets and backup trajectories is not as straightforward in other methods as it is in predictive filters, where they are calculated implicitly with the cost of solving an on-the-fly optimization.
We emphasize that the proposed centralized safety filter can be adapted for a decentralized scenario if each agent can communicate with its neighbors or predict their behavior. Developing a decentralized version of the proposed method is our next step to enhance practicality and reduce computational complexity [26]. A critical consideration when using safety filters in a non-deterministic setting is their tendency to introduce conservatism. If an accurate model of the system is unavailable, an extremely short prediction horizon and a small final set are adopted, or if there is a significant delay, safety filters may introduce unnecessary caution as any robust constrained solution. However, as the proposed modular solution is implemented as an add-on that is unaware of the cyber controller and attacks, any conservatism in the safety filter may result in the anomaly detector falsely identifying an attack. Further work is required to establish how conservatism affects anomaly detectors in the cyber layer, how the impact of conservatism on the anomaly detector can be mitigated, and whether communication between the modules may be required.